OpenStudy (anonymous):

In a word game, you choose a tile from a bag, replace it, and then choose another. If there are 28 vowels and 12 consonants, what is the probability you will choose a consonant and then a vowel?

OpenStudy (anonymous):

42/5 21/100 21/5 1/40 Listed above are all possible answers!

OpenStudy (kropot72):

The is a total of 28 + 12 = 40 tiles. The probability of choosing a consonant first is 12/40. The probability of drawing a vowel next is 28/40. The events are independent, therefore the required probability is given by: \[\frac{12}{40}\times\frac{28}{40}=you\ can\ calculate\]

OpenStudy (anonymous):

calculates to 1/1.. thats not even an option.. @kropot72

OpenStudy (kropot72):

Nope. It does calculate to one of the options. Try some simplifying by cancelling.